40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 18 02:42:28 2014, Jenne, Update, LSC, AO path partly engaged Zoom_TRXTRY4_EngagingAO_1095059162.png0020.pdf
    Reply  Thu Sep 18 17:44:55 2014, Jenne, Update, LSC, Old AO cable pulled 
       Reply  Tue Sep 30 17:26:18 2014, ericq, Update, LSC, New AO cable in place 
    Reply  Fri Sep 19 04:05:05 2014, ericq, Update, LSC, AO path partly engaged TFSR785_19-09-2014_020555.pdfTFSR785_19-09-2014_033920.pdf
       Reply  Fri Sep 19 13:12:07 2014, Jenne, Update, LSC, AO path glitches 
Message ID: 10516     Entry time: Thu Sep 18 02:42:28 2014     Reply to this: 10519   10520
Author: Jenne 
Type: Update 
Category: LSC 
Subject: AO path partly engaged 

Tonight was a night of trying to engage the AO path.  The idea was to sit at arm powers of a few on sqrtInvTrans for CARM and ALS for DARM, and try to increase the gain for REFLDC->AO path.

No exciting nit-picky details in locking procedure.  Mostly it was just a night of trying many times. 

The biggest thing that Q and I found tonight was that the 2-pin lemo cable connecting the CM board's SERVO OUT to the MC board's IN2 is shitty.  The symptom that led to this investigation was that I could increase the AO path gain arbitrarily, and have no change in the measured analog CM loop transfer function. We checked that the CM board servo out spit out signals that were roughly what we expected based on our ~2kHz excitation.  However, if we look at digitized signals from the MC board, the noise level was very high, with loads of 60Hz lines, and a teensy-tiny signal peak.  We put a small drive directly into the MC board and could see that, so we determined that the cable is bad.  We have unplugged the white 2-pin lemo, and ran a long BNC cable between the 2 boards.  Tomorrow we need to make a new 2-pin lemo cable so that we can have the lower noise differential drive signal.

After putting in the temporary cable, we do see an excitation sent to the CM board showing up after the MC board.  For this monitoring, the MC_L cable to the ADC has been borrowed, so instead of being the OUT1, the regular length signal, MC_L is currently the OUT2 monitor right after the board inputs. 

At some point in the evening, around 1:15am, ETMX started exhibiting the annoying behavior of wandering off sometimes.  I went in and pushed on the SUS cables to the satellite box, and I think it has helped, although I still saw the drift at least once after the cable-squishing.

Other than that, it has just been many trials.

The best was one where I was holding the arm powers around 4, and got the CM board's AO gain to -8 dB and the MC board's IN2 AO gain to -4 dB. I lost lock trying to increase the CM board gain to -7 dB. 

I took several transfer functions, and used Q's nifty "SRmeasure" script to gather data, and Q made a plot to see the progress.

TF progress plots:

0020.pdf

Time series of that lockloss:

Zoom_TRXTRY4_EngagingAO_1095059162.png

I don't know yet if the polarity of the CM board should be plus or minus.  This series was taken with "minus".  But,  since the phase looked opposite of Q's single arm CM board checkout from several months ago, we did a few trials with the polarity switched to "plus".  I thought we weren't getting as high of AO path gains, so I switched back to "minus", but the last few trials didn't get even as far as the plus trials did.  So, I still don't know which sign we want.

ELOG V3.1.3-